INFRARED DIODE LASER SPECTROSCOPY OF FUNDAMENTAL AND HOT BANDS OF BBr (X^{1}Σ^{+})

N. T. HUNT, W. Y. FAN, \underline{Z. LIU} and P. B. DAVIES, Department of Chemistry, Uiversity of Cambridge, Cambridge CB2 1EW, UK.

Infrared absorption spectra of ^{11}B^{79}Br, ^{11}B^{81}Br, ^{10}B^{79}Br and ^{10}B^{81}Br have been observed between 650 and 720 cm^{−1} in a BBr_{3}/He ac discharge using diode laser spectroscopy. Over 200 lines consisting of both fundamental and hot bands up to v′′ = 5 have been fitted to Dunham expansions. The fitted parameters are used to derive the vibrational constants of ^{11}B^{79}Br, ω_{e} = 685.1844(17) cm^{−1} and ω_{e}x_{e} = 3.7347(11) cm^{−1}. The equilibrium dissociation energy is also obtained from expanded Dunham-Morse potential, D_{e} = 32647(49) cm^{−1}.